OpenStudy (anonymous):

the perimeter of a rectangle is 18, and the area is 14. What is the width?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

l*w=14

OpenStudy (anonymous):

2l+2w=18

OpenStudy (anonymous):

l=7 w=2

OpenStudy (anonymous):

solve these t2

OpenStudy (anonymous):

perimeter of a rectangle=2(l+w) 2(l+w)=18 l+w=18/2 l+w=9 area of a rectangle= l*w l*w=14 so, l*w=14 and l+w=9 l=9-w l*w=14 (9-w)*w=14 9w-w^2=14 w^2-9w+14=0 w^2-7w-2w+14=0 w(w-7)-2(w-7)=0 w-2=0,w-7=0 w=2,w=7 so,length=7 and width=2